OpenStudy (anonymous):

An archer pulls back her bowstring loaded with an arrow thats .22kg. It leaves the arrow at 220 kmph, calculate work done.

OpenStudy (anonymous):

By Work Energy Theorem, Work Done = Change in Kinetic Energy KE(Initial)=0 KE(Final)=\[\frac{ mv^{2} }{ 2 }\] \[W=0-\frac{ mv^{2} }{ 2 }\] \[W=-\frac{ mv^{2} }{ 2 }\] Mass m of the bow is given, velocity v of the bow is given, make sure to convert Km/hr to m/s Use the factor of 5/18 \[v(m/s)=v(km/hr)*\frac{ 5 }{ 18 }\]

OpenStudy (anonymous):

I had it! I was using 220 instead of 61.1 though >.< Woops!